OpenStudy (anonymous):

From 200 litres of 80% concentrated milk, 30 litres is taken out and water is added. The new concentration of milk in the mixture is

OpenStudy (anonymous):

160 milk 40 water 30 liters commonly taken from both and water is added+30

OpenStudy (callisto):

Hmm.... Not really. Consider 30L of the mixture: Milk = 30 * 80% = ? Water = ?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

6 water 24 milk

OpenStudy (callisto):

Errrr, sorry. I should have asked you in this way: We take out 30L of the mixture. so we have 200 - 30 = 170L of mixture. In this 170L of mixture, how many milk and water do we have?

OpenStudy (anonymous):

136 litre milk and 34 water

OpenStudy (callisto):

Now we add 30L of water. How much water and milk do we have?

OpenStudy (anonymous):

136 and 64 so 17: 8

OpenStudy (callisto):

Done :)